# DeepC4: Deep Conditional Census-Constrained Clustering for Large-scale Multitask Spatial Disaggregation of Urban Morphology

### v1

@@ Abstract
To understand our global progress for sustainable development and disaster risk reduction in many developing economies, two recent major initiatives – the Uniform African Exposure Dataset of the Global Earthquake Model (GEM) Foundation and the Modelling Exposure through Earth Observation Routines (METEOR) Project – implemented classical spatial disaggregation techniques to generate large-scale mapping of urban morphology using the information from various satellite imagery and its derivatives, geospatial datasets of the built environment, and subnational census statistics. However, the local discrepancy with well-validated census statistics and the propagated model uncertainties remain a challenge in such coarse-to-fine-grained mapping problems, specifically constrained by weak and conditional label supervision. Therefore, we present Deep Conditional Census-Constrained Clustering (DeepC4), a novel deep learning-based spatial disaggregation approach that incorporates local census statistics as cluster-level constraints while considering multiple conditional label relationships in a joint multitask learning of the patterns of satellite imagery. To demonstrate, compared to GEM and METEOR, we enhanced the quality of Rwandan maps of urban morphology, specifically building exposure and physical vulnerability, at the third-level administrative unit from the 2022 census. As the world approaches the conclusion of our global frameworks in 2030, our work has offered a new deep learning-based mapping technique towards a spatial auditing of our existing coarse-grained derived information at large scales.
@@ 1 Introduction
As one of the key priorities of the 2015-2030 Sendai Framework for Disaster Risk Reduction (United Nations (2015), UN) alongside the targets of the 2030 Agenda for Sustainable Development (UN, 2015), the need to understand disaster risk at large scales has urged recent various multi-country projects such as the METEOR project for UN-designated Least Developed Countries (LDCs) (Huyck et al., 2019) and the Global Earthquake Model (GEM) Foundation Uniform Africa Exposure Data (Paul et al., 2022) to map the locations of populated human settlements and buildings in many developing economies. Assessing disaster risk complements this information on the exposed location with the characterization of physical vulnerability, which is commonly described in terms of the type of construction material and the height of a building. For example, the standard taxonomies of GEM (Brzev et al., 2013), PAGES-STR (Jaiswal et al., 2010), HAZUS (Kircher et al., 2006), and EMS-98 (Grünthal, 1998) provide a classification system on building typologies, indicating the inherent structural capacity or vulnerability of a building. More broadly, in a regional scale, this characterization of the physical attributes of buildings collectively defines the three-dimensional urban form or morphology of a place (Labetski et al., 2023).
@@
However, while such large-scale efforts using classical spatial disaggregation techniques have revealed a comprehensive understanding of the urban morphology to assess regional disaster risk, its numerical discrepancy with well-validated census records at the local level has remained a significant challenge in establishing a general consensus between the data-driven mapping techniques and the local knowledge and practice (Wardrop et al., 2018; Petrarulo et al., 2022), thereby limiting its use by the wider community such as urban planners and disaster risk managers on the ground (Gevaert et al., 2021). Although other finer-grained efforts such as building-level attribute identification have shown reliable predictive performances (Geiß et al., 2018; Gouveia et al., 2024; Silva et al., 2024), these alternatives are data-expensive, resource-intensive, and scope-limited, when scaling up to a larger areal extent. Thus, the tradeoff between the resolution quality of mapping and the accuracy of information has become a primary issue in supporting the evidence-based decisions of many global frameworks (Clark et al., 2023).
@@
Despite the potential wide-scale offerings of previous studies, their use of secondary products and input satellite imagery with varying temporal signatures further compounded and propagated the spatial uncertainties of the resulting maps of urban morphology. For example, even though both GEM and METEOR used the rich information from remote sensing data and the data-driven techniques such as the random forest tool (Stevens et al., 2015) for WorldPop population map (Tatem, 2017) and the symbolic machine learning (Blei et al., 2018) for Global Human Settlement Layer (Pesaresi et al., 2024), the performance of WorldPop for Kenya still achieved 60% (Stevens et al., 2015). In addition, the METEOR project used varying temporal information of their input geospatial maps due to the unavailability of temporally coherent data during the time of analysis (Huyck et al., 2019). As a result, these challenges inherent to the model uncertainties still necessitate subsequent careful validation and other independent analyses for suitability and overall quality (Gevaert et al., 2021).
@@
Therefore, we develop the Deep Conditional Census Constrained Clustering (DeepC4), a novel deep learning-based spatial disaggregation approach that fundamentally addresses the numerical discrepancy with coarse-grained or cluster-level labels (e.g., census statistics) and incorporates conditionality on multiple possible labels (i.e., from expert belief system gained from local indigenous knowledge) in a multitask optimization, while still exploiting the richness of information that can be derived from near-real-time publicly available satellite imagery and other derived products. This coarse-to-fine-grained approach can also be generalizable to other weakly supervised problems that involve the use of EO data for urban informatics applications with encoded constraints from other external validation surveys. In this study, we demonstrate the application of DeepC4 using the case of Rwanda, one of the LDCs which has experienced a significant technological difficulty in advancing exposure mapping and its physical vulnerability characteristics.
@@ 2.3 Experts’ Belief System on Construction Practice
We considered the valuable encoded mapping scheme of Paul et al. (2022), which creates discrete probabilistic relationships among wall material, height, and macro-taxonomy classes, since it is based on a comprehensive engagement and judgment of local experts on the construction methodology. This information controls the learning behavior of DeepC4 using the numerical conditions among these urban morphology indicators. Table 1 enumerates the complete list of urban morphology indicators in our resulting Rwandan maps, which were reduced from the original literature (Paul et al., 2022) due to the overlapping general characteristics and additional computational benefit in the implementation of constrained clustering (i.e., lower number of categories results in less computational demand). The discrete conditional probability values are presented in A.
@@ 2.4 Earth Observation Data
We obtained the following pre-processed 10-meter satellite imagery whose signals serve as a proxy to infer the distribution of distinct urban morphology indicators for the 2022 period via Google Earth Engine (Gorelick et al., 2017).
Sentinel-1 SAR GRD. We used the mean of the Ground Range Detected (GRD) scenes that are acquired from the dual-polarization C-band Synthetic Aperture Radar (SAR) instrument at 5.405 GHz of the Sentinel-1 satellite (Copernicus Sentinel data, 2024b). It consists of two bands: VV (vertical transmit, vertical receive) and VH (vertical transmit, horizontal receive) signals. To avoid data incompleteness across large areas, we disregarded filtering by orbital number and satellite direction.
Sentinel-2 Harmonized MSI. We also extracted the median of the atmospherically corrected surface reflectance signals represented by 10 bands – red (B4), green (B3), blue (B2), red edge 1 (B5), red edge 2 (B6), red edge 3 (B7), red edge 4 (B8A), near infrared (B8), short-wave infrared 1 (B11), and short-wave infrared 2 (B12) – that are acquired from the MultiSpectral Instrument (MSI) of the Sentinel-2 satellite (Copernicus Sentinel data, 2024a). The aggregation by year also enables minimizing the unnecessary noisy cloudy or shadowy signals using the available and corresponding Sentinel-2 cloud probability dataset (Copernicus Sentinel data, 2024c).
@@ 2.6 Building-level Groundtruth Information
While a vanilla unsupervised clustering does not necessitate any groundtruth information, our DeepC4, however, involves the training of deep neural network parameters to learn a more meaningful latent reduced representation of EO satellite imagery signals using building-level groundtruth information, which effectively supervises the constrained clustering. Because such groundtruth information in a form of a 2022 Rwandan building footprint is not widely available and may often use a different modality (e.g., different classification system), we instead used the limited supervision provided by the 2015 dataset on building typology for the Kigali City, Rwanda, which covers 30 (7.2%) out of all 416 sectors (Bachofer et al., 2019). With this, our empirical results assumed that the 2015 groundtruth information has a minimal difference with the 2022 true information by defining a subset of map locations with valid pairs of DeepC4 2022 predictions and the post-processed 2015 groundtruth information, as presented with further details in Section 3.3.5.
@@ 3.1 Preprocessing of constraints and conditions
The preprocesing started with the encoding of census statistics from 30 district-level reports into a machine-readable format. For every sector $s_{i}$, where $i=1,2,...,N_{s}$ and $N_{s}$ correspond to the total number of sectors in the country, we computed its corresponding average household size, $h_{s_{i}}$, by taking the ratio of sector-level population and private household count. Because the census also reported the composition of the population by urban and rural areas, we then computed the number of dwellings for urban ($d_{{s_{i}},urban}$) and rural ($d_{{s_{i}},rural}$) by respectively dividing the urban and rural population by $h_{s_{i}}$. By respective conditional proportions, we distributed $d_{{s_{i}},urban}$ and $d_{{s_{i}},rural}$ into wall, material, height, and macro-taxonomy classes using the probabilistic relationships in A.
As a result, for both rural and urban areas of sector $s_{i}$, we obtained the number of buildings for each urban morphology indicator in Table 1. Given a hard assumption wherein a building has a 60-m2 ground floor area similar to Paul et al. (2022), we converted the number of buildings into the number of grid pixels at 10-m scale, which is the representative spatial resolution of our EO satellite imagery signals. This follows that a 10-m grid pixel on the map, whose area is 100 m2, contains approximately 1.67 units of 60-m2 buildings (i.e., 100 m2 per grid / 60 m2 per building). We used the resulting number of grid pixels as our transformed constraints in the succeeding section.
@@ 3.3.2 Dimensional reduction
For computational efficiency and consistency of constrained clustering implementation, we extracted a low-dimension latent representation, $Z_{s_{i}}$, from the high dimensionality of $X_{s_{i}}$ using an autoencoder, a deep non-linear reduction technique. Following a fully connected layer network architecture, our study set an arbitrary number of latent channels for each urban morphology indicator, as shown in Figure 2. In this study, we used a total of three latent channels for roof ($z_{s_{i},roof}$), wall ($z_{s_{i},wall}$), and height ($z_{s_{i},height}$). Across all 20 sectors, we trained a single set of deep neural network parameters until the model performance converged, indicating that the learned latent representation has achieved a desirable clustering behavior given the conditional constraints under consideration.
@@ 3.3.3 Constrained clustering
The DeepC4 considered the conditional constraints as input to the minimum cost network flow optimization algorithm (Bradley et al., 2000) via integer linear programming (Huangfu and Hall, 2018) as an equivalent formulation. For example, if each of 1,000 grid pixels has a directed connection to every possible cluster assignment (e.g., $N$ number of roof classes), resulting in an exhaustive network with $N\times 1000$ candidate connections. In this way, the ‘flow’ signifies a selected cluster to which a particular grid pixel will be assigned. Assuming a uniform cost to trigger ‘flow’ or ‘cluster selection’ for every connection between any grid pixel and the chosen cluster, our objective is to achieve a minimum cost in such a way that the total ’flow’ to each possible cluster should follow the provided conditional constraints in the beginning because the total ’flow’ refers to the desired minimum number of grid pixels for each possible cluster assignment.
@@ 3.3.4 Joint multitask learning
It is important to note that, while the clustering is separately convex-optimizing to find its centers, it still depends on the learned latent representations as a direct result of the training of our autoencoder neural network. This joint minimization of the reconstruction loss (Equation 1) and the clustering loss (Equation 4) influencing a number of latent representations (i.e., of multiple urban morphology indicators) effectively results in a joint multitask learning. In this study, the joint total loss (Equation 5) is calculated as:
@@ 3.3.5 Performance evaluation
While clustering is an unsupervised approach, the use of deep representation learning in the form of autoencoder neural networks to infer cluster assignments still requires validation to evaluate the resulting trained autoencoder. Due to the limited sample of 20 sectors only, we performed a 5-fold cross-validation wherein we set four sectors for each fold (i.e., 20 sectors $\div$ 5 folds). For five times, we tested and averaged the performance of every model trained using only 16 remaining sectors on each unseen fold with 6 randomly selected sectors.
In addition, the overlapping characteristics of labels between Bachofer et al. (2019) ($Y_{C^{h}}$) and the target classes ($C^{h}_{s_{i}}$) present a difficulty in using standard accuracy metrics wherein the task of assigning a true negative prediction has become complicated due to the weak and limited supervision of the available building groundtruth information. Thus, we evaluated the performance of our resulting trained DeepC4 model using a modified precision (Equation 6) with multiple $OR$ operators, accounting for shared characteristics of the classes of each urban morphology indicator as presented in B.
$precision=\frac{TP_{s_{i}}}{TP_{s_{i}}+FP_{s_{i}}}=\frac{\sum X}{|M_{i}|}$ (6)
where $TP$ and $FP$ are the number of true positives and false positives, respectively, and:
@@
Equally important to the selection of the final trained model is another measure of precision, which accounts for the imbalanced number of target classes. For each sector $s_{i}$ with $k$ target classes, we used Equation 7 to compute a vector of corresponding weight factors, $r^{h}_{s_{i}}$ to be multiplied to each precision score from Equation 6.

### v2

@@ Abstract
To understand our global progress for sustainable development and disaster risk reduction in many developing economies, two recent major initiatives – the Uniform African Exposure Dataset of the Global Earthquake Model (GEM) Foundation and the Modelling Exposure through Earth Observation Routines (METEOR) Project – implemented classical spatial disaggregation techniques to generate large-scale mapping of urban morphology using the information from various satellite imagery and its derivatives, geospatial datasets of the built environment, and subnational census statistics. However, the local discrepancy with well-validated census statistics and the propagated model uncertainties remain a challenge in such coarse-to-fine-grained mapping problems, specifically constrained by weak and conditional label supervision. Therefore, we present Deep Conditional Census-Constrained Clustering (DeepC4), a novel deep learning-based spatial disaggregation approach that incorporates local census statistics as cluster-level constraints while considering multiple conditional label relationships in a joint multitask learning of the patterns of satellite imagery. To demonstrate, compared to GEM and METEOR, we enhanced the quality of Rwandan maps of urban morphology, specifically building exposure and physical vulnerability, at the third-level administrative unit from the 2022 census. As the world approaches the conclusion of many global frameworks in 2030, our work offers a new deep learning-based mapping technique that explicitly encodes well-validated census and experts’ belief systems to achieve an explainable and interpretable auditing of existing coarse-grained derived information at large scales.
@@
journal: ISPRS Journal of Photogrammetry and Remote Sensing
@@ 1 Introduction
As one of the key priorities of the 2015-2030 Sendai Framework for Disaster Risk Reduction (United Nations (2015), UN) alongside the targets of the 2030 Agenda for Sustainable Development (UN, 2015), the need to understand disaster risk at large scales has urged recent various multi-country projects such as the METEOR project for UN-designated Least Developed Countries (LDCs) (Huyck et al., 2019) and the Global Earthquake Model (GEM) Foundation Uniform Africa Exposure Data (Paul et al., 2022) to map the locations of populated human settlements and buildings in many developing economies. Assessing disaster risk complements this information on the exposed location with the characterization of physical vulnerability, which is commonly described in terms of the type of construction material and the height of a building. For example, the standard taxonomies of GEM (Brzev et al., 2013), PAGES-STR (Jaiswal et al., 2010), HAZUS (Kircher et al., 2006), and EMS-98 (Grünthal, 1998) provide a classification system on building typologies, indicating the inherent structural capacity or vulnerability of a building. More broadly, in a regional scale, this characterization of the physical attributes of buildings collectively contributes to the three-dimensional urban form or morphology of a place (Labetski et al., 2023).
@@
However, while such large-scale efforts using classical spatial disaggregation techniques have revealed a comprehensive understanding of the urban morphology indicators used for regional disaster risk assessment, its numerical discrepancy with well-validated census records at the local level has remained a significant challenge in establishing a general consensus between the data-driven mapping techniques and the local knowledge and practice (Wardrop et al., 2018; Petrarulo et al., 2022), thereby limiting its use by the wider community such as urban planners and disaster risk managers on the ground (Gevaert et al., 2021). Although other finer-grained efforts such as building-level attribute identification have shown reliable predictive performances (Geiß et al., 2018; Gouveia et al., 2024; Silva et al., 2024), these alternatives are data-expensive, resource-intensive, and scope-limited, when scaling up to a larger areal extent. Thus, the tradeoff between the resolution quality of mapping and the accuracy of information has become a primary issue in supporting the evidence-based decisions of many global frameworks (Clark et al., 2023).
@@
Despite the potential wide-scale offerings of previous studies, their use of secondary products and input satellite imagery with varying temporal signatures further compounded and propagated the spatial uncertainties of the resulting maps of urban morphology indicators. For example, even though both GEM and METEOR used the rich information from remote sensing data and the data-driven techniques such as the random forest tool (Stevens et al., 2015) for WorldPop population map (Tatem, 2017) and the symbolic machine learning (Blei et al., 2018) for Global Human Settlement Layer (Pesaresi et al., 2024), the performance of WorldPop for Kenya still achieved 60% (Stevens et al., 2015). In addition, the METEOR project used varying temporal information of their input geospatial maps due to the unavailability of temporally coherent data during the time of analysis (Huyck et al., 2019). As a result, these challenges inherent to the model uncertainties still necessitate subsequent careful validation and other independent analyses for suitability and overall quality (Gevaert et al., 2021).
@@
Therefore, we develop the Deep Conditional Census Constrained Clustering (DeepC4), a novel deep learning-based spatial disaggregation approach that fundamentally addresses the numerical discrepancy with coarse-grained or cluster-level labels (e.g., census statistics) and incorporates conditionality on multiple possible labels (i.e., from experts’ belief system gained from local indigenous knowledge) in a multitask optimization, while still exploiting the richness of information that can be derived from near-real-time publicly available satellite imagery and other derived products. This coarse-to-fine-grained approach can also be generalizable to other weakly supervised problems that involve the use of EO data for urban informatics applications with encoded constraints from other external validation surveys. By considering the explicit importance of well-validated census and experts’ belief systems, our approach also contributes to the overall consistency of resulting deep learning-based maps, thereby improving their relative explainability and interpretability with respect to the current understanding of end users. In this study, we demonstrate the application of DeepC4 using the case of Rwanda, one of the LDCs which has experienced a significant technological difficulty in advancing exposure mapping and its physical vulnerability characteristics.
@@ 2.3 Experts’ Belief System on Construction Practice
We considered the valuable encoded mapping scheme of Paul et al. (2022), which creates discrete probabilistic relationships among wall material, height, and macro-taxonomy classes, since it is based on a comprehensive engagement and judgment of local experts on the construction methodology. This information controls the learning behavior of DeepC4 using the numerical conditions among these urban morphology indicators. As presented in A, this explicit encoding of discrete conditional probability values supports the interpretability of resulting maps since the inferred distribution of classes, including any deviations, can be explained by the probabilistic relationships.
Table 1 enumerates the complete list of urban morphology indicators in our resulting Rwandan maps, which were reduced from the original literature (Paul et al., 2022) due to the overlapping general characteristics and additional computational benefit in the implementation of constrained clustering (i.e., lower number of categories results in less computational demand). However, it is important to note that, in this study, these presented indicators are limited only to physical building-level attributes, but still contribute to the understanding of the urban morphology of a place (e.g., spatial patterns, density, and other numerical metrics).
@@ 2.4 Earth Observation Data
We obtained the following pre-processed 10-meter satellite imagery whose signals serve as a proxy to infer the distribution of distinct urban morphology indicators for the 2022 period via Google Earth Engine (Gorelick et al., 2017). Although no experiment was performed to evaluate the marginal contribution of each signal band, our study relied on the inherent capabilities of each modality to capture relevant remotely sensed information and to primarily demonstrate the latent reduced representation for a deep constrained clustering implementation.
Sentinel-1 SAR GRD. Using its cloud-penetrating capability for surface roughness sensitivity (Koppel et al., 2017), we used the mean of the Ground Range Detected (GRD) scenes acquired from the dual-polarization C-band Synthetic Aperture Radar (SAR) instrument at 5.405 GHz of the Sentinel-1 satellite (Copernicus Sentinel data, 2024b). It consists of two bands: VV (vertical transmit, vertical receive) and VH (vertical transmit, horizontal receive) signals. To avoid data incompleteness across large areas, we disregarded filtering by orbital number and satellite direction.
Sentinel-2 Harmonized MSI. Using its multispectral imaging capability for optical signatures (Braun et al., 2019), we also extracted the median of the atmospherically corrected surface reflectance signals represented by 10 bands – red (B4), green (B3), blue (B2), red edge 1 (B5), red edge 2 (B6), red edge 3 (B7), red edge 4 (B8A), near infrared (B8), short-wave infrared 1 (B11), and short-wave infrared 2 (B12) – that are acquired from the MultiSpectral Instrument (MSI) of the Sentinel-2 satellite (Copernicus Sentinel data, 2024a). The aggregation by year also enables minimizing the unnecessary noisy cloudy or shadowy signals using the available and corresponding Sentinel-2 cloud probability dataset (Copernicus Sentinel data, 2024c).
@@ 2.6 Building-level Groundtruth Information
While a vanilla unsupervised clustering does not necessitate any groundtruth information, our DeepC4, however, involves the training of deep neural network parameters to learn a more meaningful latent reduced representation of EO satellite imagery signals using building-level groundtruth information, which effectively supervises the constrained clustering. Because such groundtruth information in a form of a 2022 Rwandan building footprint is not widely available and may often use a different modality (e.g., different classification system), we instead used the weak and limited supervision provided by the 2015 dataset on building typology for the Kigali City, Rwanda, which covers 30 (7.2%) out of all 416 sectors (Bachofer et al., 2019).
As a limitation, our empirical results assumed that the 2015 groundtruth information has a minimal difference with the 2022 true information by defining a common subset of map locations with valid pairs of the post-processed 2015 groundtruth information and DeepC4 2022 predictions, as presented with further details in Section 3.3.5. This implies that any subsequent interpretations are contextually limited by the geographic and temporal biases inherent in using only the available supervising information from Kigali City, Rwanda, thereby potentially disregarding distinct urban-rural differences. Nonetheless, while it may be desirable to have representative samples across the country of Rwanda, our demonstration recognizes that this current data gap reflects an important domain challenge and presents an opportunity for future fine-tuning improvement, when a large-scale survey of Rwandan building typology becomes available.
@@ 3.1 Preprocessing of constraints and conditions
The preprocesing started with the encoding of census statistics from 30 district-level reports into a machine-readable format. For every sector $s_{i}$, where $i=1,2,...,N_{s}$ and $N_{s}$ correspond to the total number of sectors in the country, we computed its corresponding average household size, $h_{s_{i}}$, by taking the ratio of sector-level population and private household count. Because the census also reported the composition of the population by urban and rural areas, we then computed the number of dwellings for urban ($d_{{s_{i}},urban}$) and rural ($d_{{s_{i}},rural}$) by respectively dividing the urban and rural population by $h_{s_{i}}$. By respective conditional proportions, we distributed $d_{{s_{i}},urban}$ and $d_{{s_{i}},rural}$ into wall, material, height, and macro-taxonomy classes using the probabilistic relationships in A. As a result, for both rural and urban areas of sector $s_{i}$, we obtained the number of buildings for each urban morphology indicator in Table 1.
However, due to the lack of temporally consistent data on building footprints in a vectorized or polygonal format, as previously discussed in Section 2.5, we transformed the building counts by approximating their equivalents to the rasterized, image, or pixel-based representation. We note that this approach may be sensitive to the vector-to-raster conversion because of the differences between the grid configuration and the orientation and geometry of the enclosed building footprints.
Thus, to enable a coherent rasterized analysis, by using a hard assumption wherein a building has a 60-m2 ground floor area similar to Paul et al. (2022), we converted the number of buildings into the number of grid pixels at 10-m scale, which is the representative spatial resolution of our EO satellite imagery signals. This follows that a 10-m grid pixel on the map, whose area is 100 m2, contains approximately 1.67 units of 60-m2 buildings (i.e., 100 m2 per grid / 60 m2 per building). We used the resulting number of grid pixels as our transformed constraints in the succeeding section. When high-resolution imagery becomes available, it follows that this approximating assumption can remain valid, with further reduced uncertainty brought about by the sensitivity to vector-to-raster conversion.
@@ 3.3.2 Dimensional reduction
For computational efficiency and consistency of constrained clustering implementation, we extracted a low-dimension latent representation, $Z_{s_{i}}$, from the high dimensionality of $X_{s_{i}}$ using an autoencoder, a deep non-linear reduction technique. While the use of other image-based deep learning techniques may be desirable to capture spatial or positional information, these approaches require an approximate transformation of cluster-level constraints into corresponding image tiles. The autoencoder addressed this limitation by incorporating $X_{s_{i}}$ with location encoding features in a tabular format, as the cluster-level constraints are only expressed at the sector level and not in a form suitable for image tile representation.
Following a fully connected layer network architecture, our study assigned an arbitrary number of latent channels to each urban morphology indicator, as shown in Figure 2. In this study, we used a total of three latent channels for roof ($z_{s_{i},roof}$), wall ($z_{s_{i},wall}$), and height ($z_{s_{i},height}$) in such a way that each separate latent channel is correspondingly interpretable to individual urban morphology indicator. Across all 20 sectors, we trained a single set of deep neural network parameters until the model performance converged, indicating that the learned latent representation has achieved a desirable clustering behavior given the conditional constraints under consideration.
@@ 3.3.3 Constrained clustering
The DeepC4 considered the conditional constraints as input to the minimum cost network flow optimization algorithm (Bradley et al., 2000) via integer linear programming (Huangfu and Hall, 2018) as an equivalent formulation. For example, if each of 1,000 grid pixels has a directed connection to every possible cluster assignment (e.g., $N$ number of roof classes), resulting in an exhaustive network with $N\times 1000$ candidate connections. In this way, the ‘flow’ signifies a selected cluster to which a particular grid pixel will be assigned. Assuming a uniform cost to trigger ‘flow’ or ‘cluster selection’ for every connection between any grid pixel and the chosen cluster, our objective is to achieve a minimum cost in such a way that the total ’flow’ to each possible cluster should follow the provided conditional constraints in the beginning because the total ’flow’ refers to the desired minimum number of grid pixels for each possible cluster assignment, as presented in Table 1.
@@ 3.3.4 Joint multitask learning
It is important to note that, while the clustering is separately convex-optimizing to find its centers, it still depends on the learned latent representations as a direct result of the training of our autoencoder neural network. This joint minimization of the reconstruction loss (Equation 1) and the clustering loss (Equation 4) influencing a number of latent representations (i.e., of multiple urban morphology indicators) effectively results in a joint multitask learning. In this study, we calculated the joint total loss (Equation 5) as unweighted because our initial training implementation suggested comparable orders of loss magnitude while considering the differences in the discreteness of latent labels and the continuous nature of reconstructed signals.
@@ 3.3.5 Performance evaluation
While clustering is an unsupervised approach, the use of deep representation learning in the form of autoencoder neural networks to infer cluster assignments still requires validation to evaluate the resulting trained autoencoder. Due to the limited sample of 20 sectors only, we performed a five-fold cross-validation wherein we set four sectors for each fold (i.e., 20 sectors $\div$ 5 folds). For five times, we tested and averaged the performance of every model trained using only 16 remaining sectors on each unseen fold with six randomly selected sectors.
In addition, the overlapping characteristics of labels (i.e., many-to-many mapping) between Bachofer et al. (2019) ($Y_{C^{h}}$) and the target classes ($C^{h}_{s_{i}}$) present a difficulty in using standard accuracy metrics, accounting for shared characteristics of the classes of each urban morphology indicator. The task of assigning binary classification metrics has become complicated (i.e., ambiguous or ill-defined negatives) due to the weak and limited supervision of the available building groundtruth information in a multiclass context. Nonetheless, the validity of predictions can still be conservatively verified against the three derived mapping tables in B. Thus, we reported the performance of our resulting trained DeepC4 model using the proportion of valid predictions (Equation 6) with multiple $OR$ operators, which measures the fraction of predictions that satisfy these limitations of label ambiguity. This simplifies the performance evaluation of whether the trained DeepC4 model results in a valid (i.e., with a check symbol in the mapping table) or invalid (i.e., otherwise) prediction.
$p_{valid}=\frac{\sum X}{|M_{i}|}$ (6)
where $|M_{i}|$ indicate the number of grid pixels and:
@@
Equally important to the selection of the final trained model is a modified $p_{valid}$, which accounts for the imbalanced number of target classes. For each sector $s_{i}$ with $k$ target classes, we used Equation 7 to compute a vector of corresponding weight factors, $r^{h}_{s_{i}}$ to be multiplied to each $p_{valid}$ from Equation 6.